OpenStudy (anonymous):

solve: 1/2k-3=2-3/4k My answer i got was -20 (i don't think its correct)

OpenStudy (superdavesuper):

Very close but no, -20 is not right. Start by getting the terms with k on one side and everything else on the other side of = wat wuld u get?

OpenStudy (anonymous):

-3=2 - 1 and 1/4k

OpenStudy (superdavesuper):

should be 1/2k + 3/4k = 2 + 3 right?

OpenStudy (anonymous):

ok yes

OpenStudy (superdavesuper):

heheee then now what? :P

OpenStudy (anonymous):

i don't know add 2+3 which is 5?

OpenStudy (superdavesuper):

And how about the other side - add those together plz?

OpenStudy (anonymous):

1 and 1/4

OpenStudy (superdavesuper):

5/4k = 5 so k=?

OpenStudy (anonymous):

wait what?

OpenStudy (superdavesuper):

1 and 1/4 = 5/4 ok?

OpenStudy (anonymous):

well how did u get that...

OpenStudy (superdavesuper):

Sorry I should go back: 1/2k + 3/4k right So it is the same as 2/4k + 3/4k good so far?

OpenStudy (anonymous):

ok

OpenStudy (superdavesuper):

2/4k+3/4k=5/4k dats how i got it :)

OpenStudy (anonymous):

ok

OpenStudy (superdavesuper):

5/4k=5 so k=.....